OpenStudy (anonymous):

A baseball moving at 20m/s is struck by a bat and moves to the opposite direction at 30m/s. If the impact lasted 0.010s, what was the baseball's acceleration? (Please tell me what formula should I use to solve this)

OpenStudy (amistre64):

prolly a combination of a few of them

OpenStudy (amistre64):

momentum formulas might be useful

OpenStudy (amistre64):

we dont know force, and we dont know mass ... so impulse may not be useful

OpenStudy (anonymous):

how do we solve for that?

OpenStudy (amistre64):

i think i got it ...

OpenStudy (amistre64):

\[F=ma;~a=\frac Fm\] \[I=F~\Delta t=m~\Delta v;~\frac{F}{m}=\frac{\Delta v}{\Delta t }\]

OpenStudy (anonymous):

@,@ i dont get it. really. hahaha sorry. i just dont understand physics. we have our mid quarter exams tomorrow in physics and i dont know how to solve for this stuffs

OpenStudy (amistre64):

these formulas would have had to have already been presented to you thru the course of your term

OpenStudy (amistre64):

Newtons second law is: Force = mass times acceleration; and mathing that out we get a=F/m

OpenStudy (anonymous):

im in my senior year in high school. from philippines

OpenStudy (amistre64):

Impulse, change in momentum is the formula: Force, times change in time = mass, times change in velocity

OpenStudy (anonymous):

our just started last month. i believe this lesson is about vectors.

OpenStudy (amistre64):

it includes vectors yes ... but they are incident to it

OpenStudy (amistre64):

the change in velocity: 20m/s - - 30m/s = 50 m/s the change in times is given as .01 seconds 50m/s/.01s = 5000m/s^2

OpenStudy (anonymous):

so, thats the answer?

OpenStudy (amistre64):

thats the answer i come up with ... if its the correct answer or not might still be up for debate